# Ravan: Multi-Head Low-Rank Adaptation for Federated Fine-Tuning

### v1

@@ 1 Introduction
However, LoRA-based methods are highly affected by client data heterogeneity (see Table 1) because restricting updates to a low-rank subspace deprives the model of the capacity needed to fit the diverse directions introduced by heterogeneous data. Moreover, directly extending LoRA to FL, as done in FedIT [51], leads to an inexactness problem during aggregation. Since BA is a proxy for the true model update $\Delta\textbf{W}$, averaging the B and A parameters separately would be inconsistent with the true model update:
@@
where $\mathcal{C}^{(t)}$ is the selected client set at round $t$. Previous works that seek to address this exact aggregation issue suffer from accuracy loss and poor scalability in practical FL settings due to data and computational heterogeneity. FFA-LoRA [43] manages exact updates by freezing the A parameter at initialization but reduces the model expressivity relative to vanilla LoRA. FedEx-LoRA [41] adds the inexact residual $\frac{1}{|\mathcal{C}^{(t)}|}\sum_{c\in\mathcal{C}^{(t)}}\textbf{B}_{c}^{(t)}\textbf{A}_{c}^{(t)}-(\frac{1}{|\mathcal{C}^{(t)}|}\sum_{c\in\mathcal{C}^{(t)}}\textbf{B}_{c}^{(t)})(\frac{1}{|\mathcal{C}^{(t)}|}\sum_{c\in\mathcal{C}^{(t)}}\textbf{A}_{c}^{(t)})$ to the original pretrained weights W to get an exact update every round. However, the method substantially increases the communication cost of fine-tuning since the updated model weights also have to be communicated every round. Critically, these LoRA-based methods can afford only small ranks within a limited parameter budget. When the true update is high-rank, this approximation disregards most of the update’s variance and limits accuracy. Fed-SB [42], motivated by the update approximation from LoRA-XS [5], introduces a third LoRA parameter between the standard B and A parameters and only fine-tunes this additional parameter. However, it necessitates an initial round of full-parameter fine-tuning to initialize B and A, which is prohibitively expensive in FL. Furthermore, the initialization may become stale as training progresses due to data heterogeneity and partial participation.
@@ 3 Proposed Method
[…]
The pretrained W along with each $\textbf{B}_{i}$ and $\textbf{A}_{i}$ are frozen at the start of training. As a consequence, only the $\textbf{H}_{i}$ parameters and the lightweight scaling factors $s_{i}$ are updated and communicated throughout the FL training procedure. The pseudocode of the proposed method Ravan is given in Algorithm 1, and the following sections highlight key components of our framework. Specifically, we examine the importance of initialization in improving the update approximation (Section 3.1). We additionally analyze strategies for per-client head subset selection and aggregation for computational heterogeneity (Section 3.2). Together, these design choices let Ravan match the communication cost of vanilla LoRA, while delivering higher-rank, resource-aware updates that preserve exact aggregation.
@@ 3.1 Parameter Initialization
3: Init$(\textbf{B}_{i},\textbf{H}_{i}^{(0)},\textbf{A}_{i})$, $i=[1,\ldots,h]$ for $\mathcal{M}$
@@
11:        $\mathcal{H}_{c}^{(t)}\leftarrow$ SelectHeads$(c,t)$
@@ 3.2 Head Selection Strategies
  • Random Scoring: $\rho_{c,i}^{(t)}\sim\text{Unif}(0,1)$. Heads receive random scores, so clients form their fine-tuning subset by uniformly sampling heads at random.
@@
$\ i\ \text{is among the top $K_{c}$ heads}\}$ where the value of $K_{c}$ depends on client $c$’s computational constraints. During local fine-tuning, client $c$ only updates heads in $\mathcal{H}_{c}^{(t)}$. Let $\mathcal{C}_{i}^{(t)}$ denote the set of clients that fine-tuned head $i$ in communication round $t$. The server performs the update:
@@ 4.1 Experimental Setup
Table 2: Performance comparison on CIFAR-100 and SVHN.
CIFAR-100 (Acc. %) SVHN (Acc. %)
@@ 20 Clients 50 Clients 20 Clients 50 Clients
Method Rank I.I.D. Non-I.I.D. I.I.D. Non-I.I.D. I.I.D. Non-I.I.D. I.I.D. Non-I.I.D.
Full-FT N/A $89.89$ $86.86$ $89.78$ $85.17$ $95.06$ $90.29$ $94.90$ $89.49$
[…]
FedEx-LoRA $32$ $80.56$ $67.45$ $77.82$ $66.58$ $91.94$ $84.30$ $91.51$ $81.63$
FFA-LoRA $64$ $78.82$ $56.34$ $78.17$ $59.98$ $91.53$ $86.03$ $91.82$ $83.30$
Fed-SB $221$ $79.27$ $71.48$ $79.06$ $69.51$ $90.94$ $82.25$ $92.74$ $85.30$
@@
[…]
FedEx-LoRA $64$ $79.38$ $50.47$ $79.42$ $57.86$ $91.16$ $74.04$ $92.01$ $74.84$
FFA-LoRA $128$ $81.39$ $70.31$ $82.13$ $66.81$ $91.95$ $88.06$ $92.07$ $84.24$
Fed-SB $313$ $83.03$ $73.12$ $83.90$ $71.13$ $92.29$ $86.89$ $92.78$ $82.46$
@@
Method Rank I.I.D. Non-I.I.D. I.I.D. Non-I.I.D. I.I.D. Non-I.I.D. I.I.D. Non-I.I.D.
Full-FT N/A $71.34$ $69.29$ $71.71$ $70.13$ $62.19$ $62.25$ $62.41$ $62.51$
[…]
FedEx-LoRA $32$ $69.04$ $62.52$ $\mathbf{68.19}$ $63.33$ $60.99$ $\mathbf{60.68}$ $\mathbf{61.40}$ $60.56$
FFA-LoRA $64$ $68.11$ $62.36$ $68.00$ $64.86$ $60.31$ $60.40$ $61.21$ $60.14$
Fed-SB $221$ $67.15$ $63.10$ $66.69$ $63.98$ $59.93$ $59.73$ $59.96$ $60.01$
@@
[…]
FedEx-LoRA $64$ $68.59$ $65.11$ $67.75$ $64.31$ $61.23$ $60.36$ $61.43$ $60.06$
FFA-LoRA $128$ $69.33$ $66.22$ $68.42$ $64.86$ $61.50$ $60.50$ $61.66$ $60.12$
Fed-SB $313$ $68.07$ $64.18$ $67.58$ $65.59$ $60.22$ $60.11$ $60.28$ $60.60$
@@ 4.3 Ablation Studies and Analysis
Table 4 compares the initializations for the fixed bases $\textbf{B}_{i},\,\textbf{A}_{i}$ as described in Section 3.1. The fully orthogonal Gram-Schmidt initialization is consistently best on the vision tasks, adding 1.2% over random normal on SVHN and outperforming the constant baseline by 20% on CIFAR-100. On language tasks, the random normal initialization outperforms the constant baseline by 8.9% on 20 Newsgroups and the shared subspace baseline by 0.26% on MRQA. Thus, the proposed orthogonal initializations maximize the effective rank of the update and yield the strongest accuracy across all domains. While Gram-Schmidt outperforms other schemes on the vision tasks, the procedure is more computationally expensive in high-dimensional space. However, since this initialization is a one-time server-run operation at the start of training, amortized across the entire FL procedure, it adds virtually no overhead to the fine-tuning workload.
@@
We emulate devices with unequal trainable parameter budgets by drawing each client’s trainable parameter budget from three fixed distributions (bell-shaped, uniform, and skewed right). Details on the individual distributions can be found in the Appendix. As displayed in Figure 3, in these settings, all Ravan variants outperform the rank-adaptive baselines HetLoRA and FlexLoRA on both CIFAR-100 and SVHN. On CIFAR-100, the strongest baseline loses 11% of overall accuracy when moving from the bell-shaped distribution to the skewed right distribution. In comparison, Ravan only loses 2% on average across its 3 variants. A key reason is that FlexLoRA performs a per-client SVD and HetLoRA performs hard-rank truncation when redistributing the global model to individual clients. Ravan avoids these approximation operations, so its updates remain accurate even in the extreme skewed right case. We note, however, that weight-based scoring consistently lags behind the other two scoring mechanisms because it always selects the same high-magnitude heads across all clients. Random scoring and gradient-based scoring more evenly distribute updates across all heads and are better suited for heterogeneous fine-tuning in FL.
@@
Method MNLI-MM MNLI-M QNLI QQP SST-2 RTE Average
@@
FedEx-LoRA $84.15$ $84.70$ $82.74$ $86.07$ $94.61$ $65.34$ $82.94$
FFA-LoRA $85.05$ $\mathbf{85.78}$ $82.07$ $84.40$ $94.38$ $62.46$ $82.36$
Fed-SB $84.88$ $85.23$ $82.84$ $84.23$ $94.95$ $\mathbf{67.15}$ $83.21$
@@
We demonstrate the scalability of Ravan for larger model architectures by benchmarking the method against prior baselines on the GLUE benchmark [46] using LLaMA3.2-1B [17] (see Table 6). For each subtask, we use $\mathcal{C}=20$ clients and follow the training procedure described in Section 4.2. We use a trainable parameter budget of $N_{\text{total}}=3.5$ M, which corresponds to a vanilla LoRA rank of 32 as in the upper half of Tables 2 and 3. On average, Ravan outperforms the next best baseline by 0.7% with the maximum gain on a single subtask being the 1.2% improvement over Fed-SB on the QNLI dataset. This demonstrates that Ravan scales smoothly from 85 M to billion-parameter LLMs and complements state-of-the-art on-device models.
@@ A.1 Related Works
Federated learning (FL) enables a distributed set of clients to collaboratively train a single global model using their local data [35]. The primary challenges in FL are data heterogeneity and computational heterogeneity induced by the statistical and hardware differences across clients, respectively. Several works aim to address these challenges by altering either the client training procedure or the server aggregation algorithm. FedProx [31], SCAFFOLD [27], and FedDyn [1] add a corrective regularization term to each client’s local objective—whether a proximal penalty (FedProx), control-variate correction (SCAFFOLD), or dynamic regularizer (FedDyn)—to keep local updates from drifting too far from the global model, a phenomena more often referred to as “client drift”. An alternative approach is to adjust the server aggregation scheme to reduce client drift. For example, FedVARP [24] incorporates historical client updates into the current round’s aggregation step to reduce the variance caused by partial client participation and heterogeneity. FedExP [25] varies the server learning rate by using an extrapolation rule that increases the server step size when consecutive aggregated updates point in similar directions and shrinks it when they diverge.
@@
Recently, many works have adopted the pretrain-then-fine-tune framework, in which a general-purpose large language model (LLM) is adapted to a smaller downstream task [15, 40, 13, 21]. The excessive computational cost of fine-tuning LLMs has led to parameter-efficient fine-tuning (PEFT) methods that fine-tune a fraction of the overall parameters of the model. Adapter tuning [20], BitFit [6], and low-rank adaptation (LoRA) [22] have emerged as effective PEFT methods that can significantly reduce the number of parameters while maintaining task performance. Since its conception, many works have improved upon the initial LoRA formulation in various ways. Works such as QLoRA [12] and LoftQ [32] quantize the pretrained model weights to further improve the memory efficiency of LoRA-based fine-tuning. LoRA-XS [5], LoRA-SB [39], and MoRA [26] introduce an additional LoRA parameter while freezing both the model backbone and the original LoRA parameters during fine-tuning. HydraLoRA [44], LoRAMoE [16], and MoLE [48] employ a mixture-of-experts architecture to traditional LoRA-based fine-tuning frameworks.
@@
As newer applications look to use on-device data to fine-tune LLMs, PEFT methods for FL have become increasingly relevant. FedPETuning [53], FedPrompt [54], and FedIT [51] incorporate adapter tuning, prompt tuning, and LoRA into federated frameworks, respectively. More recently, methods like FFA-LoRA [43], FedEx-LoRA [41], Fed-SB [42], and RoLoRA [10] optimize LoRA in homogeneous-compute FL by addressing the inexactness in LoRA aggregation caused by separately averaging the B and A parameters. An orthogonal direction is explored by works such as HetLoRA [11], FlexLoRA [4], and FLoRA [47] that address computational heterogeneity in LoRA-based FL fine-tuning by allowing clients to train LoRA parameters with different ranks. However, an examination of cross-device fine-tuning remains limited in this context as methods like FLoRA have communication costs that scale linearly with the number of clients and communication rounds, making fine-tuning particularly difficult in large-scale FL systems. The goal of Ravan is to design a PEFT method for FL that addresses data and computational heterogeneity, while scaling effectively to cross-device settings with a large number of clients and communication rounds. In this way, we overcome shortcomings in prior works and enable resource-aware fine-tuning.

### v2

@@ 1  Introduction
However, LoRA-based methods are highly affected by client data heterogeneity (see Table 1) because restricting updates to a low‑rank subspace deprives the model of the capacity needed to fit the diverse directions introduced by heterogeneous data. Moreover, directly extending LoRA to FL, as done in FedIT [50], leads to an inexactness problem during aggregation. Since BA is a proxy for the true model update $\Delta\textbf{W}$, averaging the B and A parameters separately would be inconsistent with the true model update:
@@
where $\mathcal{C}^{(t)}$ is the selected client set at round $t$. Previous works that seek to address this exact aggregation issue suffer from accuracy loss and poor scalability in practical FL settings due to data and computational heterogeneity. FFA-LoRA [42] manages exact updates by freezing the A parameter at initialization but reduces the model expressivity relative to vanilla LoRA. FedEx-LoRA [41] adds the inexact residual $\frac{1}{|\mathcal{C}^{(t)}|}\sum_{c\in\mathcal{C}^{(t)}}\textbf{B}_{c}^{(t)}\textbf{A}_{c}^{(t)}-(\frac{1}{|\mathcal{C}^{(t)}|}\sum_{c\in\mathcal{C}^{(t)}}\textbf{B}_{c}^{(t)})(\frac{1}{|\mathcal{C}^{(t)}|}\sum_{c\in\mathcal{C}^{(t)}}\textbf{A}_{c}^{(t)})$ to the original pretrained weights W to get an exact update every round. However, the method substantially increases the communication cost of fine-tuning since the updated model weights also have to be communicated every round. Critically, these LoRA-based methods can afford only small ranks within a limited parameter budget. When the true update is high‑rank, this approximation disregards most of the update’s variance and limits accuracy. Fed-SB [40], motivated by the update approximation from LoRA-XS [bałazy2024loraxslowrankadaptationextremely], introduces a third LoRA parameter between the standard B and A parameters and only fine-tunes this additional parameter. However, it necessitates an initial round of full-parameter fine-tuning to initialize B and A, which is prohibitively expensive in FL. Furthermore, the initialization may become stale as training progresses due to data heterogeneity and partial participation.
@@ 3  Proposed Method
$\underbrace{\left(\textbf{W}\;+\;\sum\limits_{i=1}^{h}s_{i}\textbf{B}_{i}\textbf{H}_{i}\textbf{A}_{i}\right)\boldsymbol{x}}_{\textsc{Ravan}\text{ Forward Pass}}$ (7)
The pretrained W along with each $\textbf{B}_{i}$ and $\textbf{A}_{i}$ are frozen at the start of training. As a consequence, only the $\textbf{H}_{i}$ parameters and the lightweight scaling factors $s_{i}$ are updated and communicated throughout the FL training procedure. The pseudocode of the proposed method Ravan is given in Algorithm 1, and the following sections highlight key components of our framework. Specifically, we examine the importance of initialization in improving the update approximation (Section 1). We additionally analyze strategies for per-client head subset selection and aggregation for computational heterogeneity (Section 3.2). Together, these design choices let Ravan match the communication cost of vanilla LoRA, while delivering higher-rank, resource-aware updates that preserve exact aggregation.
@@ 3.1  Parameter Initialization
[…]
@@
11:    $\mathcal{H}_{c}^{(t)}\leftarrow$ SelectHeads(c,t)(c,t)
@@ 3.2  Head Selection Strategies
  • Random Scoring: $\rho_{c,i}^{(t)}\sim\text{Unif}(0,1)$. Heads receive random scores, so clients form their fine‑tuning subset by uniformly sampling heads at random.
@@
[…]
@@ 4.1  Experimental Setup
Table 2: Performance comparison on CIFAR‑100 and SVHN.
CIFAR‑100 (Acc. %) SVHN (Acc. %)
@@ 20 Clients 50 Clients 20 Clients 50 Clients
Method Rank I.I.D. Non‑I.I.D. I.I.D. Non‑I.I.D. I.I.D. Non‑I.I.D. I.I.D. Non‑I.I.D.
Full‑FT N/A $89.89$ $86.86$ $89.78$ $85.17$ $95.06$ $90.29$ $94.90$ $89.49$
[1pt/2pt] $N_{\text{total}}=1.2\,\text{M}$ FedIT $32$ $83.49$ $68.66$ $81.75$ $68.15$ $88.66$ $84.00$ $91.67$ $77.53$
FedEx‑LoRA $32$ $80.56$ $67.45$ $77.82$ $66.58$ $91.94$ $84.30$ $91.51$ $81.63$
FFA‑LoRA $64$ $78.82$ $56.34$ $78.17$ $59.98$ $91.53$ $86.03$ $91.82$ $83.30$
Fed‑SB $221$ $79.27$ $71.48$ $79.06$ $69.51$ $90.94$ $82.25$ $92.74$ $85.30$
@@
[1pt/2pt] $N_{\text{total}}=2.4\,\text{M}$ FedIT $64$ $83.82$ $71.01$ $84.04$ $73.23$ $91.39$ $84.68$ $92.06$ $79.31$
FedEx‑LoRA $64$ $79.38$ $50.47$ $79.42$ $57.86$ $91.16$ $74.04$ $92.01$ $74.84$
FFA‑LoRA $128$ $81.39$ $70.31$ $82.13$ $66.81$ $91.95$ $88.06$ $92.07$ $84.24$
Fed‑SB $313$ $83.03$ $73.12$ $83.90$ $71.13$ $92.29$ $86.89$ $92.78$ $82.46$
@@
Method Rank I.I.D. Non‑I.I.D. I.I.D. Non‑I.I.D. I.I.D. Non‑I.I.D. I.I.D. Non‑I.I.D.
Full‑FT N/A $71.34$ $69.29$ $71.71$ $70.13$ $62.19$ $62.25$ $62.41$ $62.51$
[1pt/2pt] $N_{\text{total}}=2.4\,\text{M}$ FedIT $32$ $\mathbf{69.07}$ $61.98$ $67.99$ $60.67$ $61.00$ $60.57$ $61.24$ $60.52$
FedEx‑LoRA $32$ $69.04$ $62.52$ $\mathbf{68.19}$ $63.33$ $60.99$ $\mathbf{60.68}$ $\mathbf{61.40}$ $60.56$
FFA‑LoRA $64$ $68.11$ $62.36$ $68.00$ $64.86$ $60.31$ $60.40$ $61.21$ $60.14$
Fed‑SB $221$ $67.15$ $63.10$ $66.69$ $63.98$ $59.93$ $59.73$ $59.96$ $60.01$
@@
[1pt/2pt] $N_{\text{total}}=4.7\,\text{M}$ FedIT $64$ $\mathbf{69.36}$ $64.41$ $68.12$ $62.67$ $61.25$ $60.75$ $61.39$ $60.26$
FedEx‑LoRA $64$ $68.59$ $65.11$ $67.75$ $64.31$ $61.23$ $60.36$ $61.43$ $60.06$
FFA‑LoRA $128$ $69.33$ $66.22$ $68.42$ $64.86$ $61.50$ $60.50$ $61.66$ $60.12$
Fed‑SB $313$ $68.07$ $64.18$ $67.58$ $65.59$ $60.22$ $60.11$ $60.28$ $60.60$
@@ 4.3  Ablation Studies and Analysis
Table 4 compares the initializations for the fixed bases $\textbf{B}_{i},\,\textbf{A}_{i}$ as described in Section 1. The fully orthogonal Gram-Schmidt initialization is consistently best on the vision tasks, adding 1.2% over random normal on SVHN and outperforming the constant baseline by 20% on CIFAR-100. On language tasks, the random normal initialization outperforms the constant baseline by 8.9% on 20 Newsgroups and the shared subspace baseline by 0.26% on MRQA. Thus, the proposed orthogonal initializations maximize the effective rank of the update and yield the strongest accuracy across all domains. While Gram-Schmidt outperforms other schemes on the vision tasks, the procedure is more computationally expensive in high-dimensional space. However, since this initialization is a one-time server-run operation at the start of training, amortized across the entire FL procedure, it adds virtually no overhead to the fine-tuning workload.
@@
We emulate devices with unequal trainable parameter budgets by drawing each client’s trainable parameter budget from three fixed distributions (bell-shaped, uniform, and skewed right). Details on the individual distributions can be found in the Appendix. As displayed in Figure 3, in these settings, all Ravan variants outperform the rank-adaptive baselines HetLoRA and FlexLoRA on both CIFAR-100 and SVHN. On CIFAR-100, the strongest baseline loses 11% of overall accuracy when moving from the bell-shaped distribution to the skewed right distribution. In comparison, Ravan only loses 2% on average across its 3 variants. A key reason is that FlexLoRA performs a per‑client SVD and HetLoRA performs hard‑rank truncation when redistributing the global model to individual clients. Ravan avoids these approximation operations, so its updates remain accurate even in the extreme skewed right case. We note, however, that weight-based scoring consistently lags behind the other two scoring mechanisms because it always selects the same high‑magnitude heads across all clients. Random scoring and gradient-based scoring more evenly distribute updates across all heads and are better suited for heterogeneous fine-tuning in FL.
@@
Method MNLI‑$M$ MNLI‑M QNLI QQP SST‑2 RTE Average
@@
FedEx‑LoRA $84.15$ $84.70$ $82.74$ $86.07$ $94.61$ $65.34$ $82.94$
FFA‑LoRA $85.05$ $\mathbf{85.78}$ $82.07$ $84.40$ $94.38$ $62.46$ $82.36$
Fed‑SB $84.88$ $85.23$ $82.84$ $84.23$ $94.95$ $\mathbf{67.15}$ $83.21$
@@
We demonstrate the scalability of Ravan for larger model architectures by benchmarking the method against prior baselines on the GLUE benchmark [45] using LLaMA3.2-1B [16] (see Table 6). For each subtask, we use $\mathcal{C}=20$ clients and follow the training procedure described in Section 4.2. We use a trainable parameter budget of $N_{\text{total}}=3.5$ M, which corresponds to a vanilla LoRA rank of 32 as in the upper half of Tables 2 and 3. On average, Ravan outperforms the next best baseline by 0.7% with the maximum gain on a single subtask being the 1.2% improvement over Fed-SB on the QNLI dataset. This demonstrates that Ravan scales smoothly from 85 M to billion‑parameter LLMs and complements state-of-the-art on-device models.
@@ A.1  Related Works
Federated learning (FL) enables a distributed set of clients to collaboratively train a single global model using their local data [34]. The primary challenges in FL are data heterogeneity and computational heterogeneity induced by the statistical and hardware differences across clients, respectively. Several works aim to address these challenges by altering either the client training procedure or the server aggregation algorithm. FedProx [30], SCAFFOLD [26], and FedDyn [1] add a corrective regularization term to each client’s local objective—whether a proximal penalty (FedProx), control‑variate correction (SCAFFOLD), or dynamic regularizer (FedDyn)—to keep local updates from drifting too far from the global model, a phenomena more often referred to as “client drift”. An alternative approach is to adjust the server aggregation scheme to reduce client drift. For example, FedVARP [23] incorporates historical client updates into the current round’s aggregation step to reduce the variance caused by partial client participation and heterogeneity. FedExP [24] varies the server learning rate by using an extrapolation rule that increases the server step size when consecutive aggregated updates point in similar directions and shrinks it when they diverge.
@@
Recently, many works have adopted the pretrain-then-fine-tune framework, in which a general-purpose large language model (LLM) is adapted to a smaller downstream task [14, 39, 12, 20]. The excessive computational cost of fine-tuning LLMs has led to parameter-efficient fine-tuning (PEFT) methods that fine-tune a fraction of the overall parameters of the model. Adapter tuning [19], BitFit [5], and low-rank adaptation (LoRA) [21] have emerged as effective PEFT methods that can significantly reduce the number of parameters while maintaining task performance. Since its conception, many works have improved upon the initial LoRA formulation in various ways. Works such as QLoRA [11] and LoftQ [31] quantize the pretrained model weights to further improve the memory efficiency of LoRA-based fine-tuning. LoRA-XS [bałazy2024loraxslowrankadaptationextremely], LoRA-SB [38], and MoRA [25] introduce an additional LoRA parameter while freezing both the model backbone and the original LoRA parameters during fine-tuning. HydraLoRA [43], LoRAMoE [15], and MoLE [47] employ a mixture-of-experts architecture to traditional LoRA-based fine-tuning frameworks.
@@
As newer applications look to use on-device data to fine-tune LLMs, PEFT methods for FL have become increasingly relevant. FedPETuning [52], FedPrompt [53], and FedIT [50] incorporate adapter tuning, prompt tuning, and LoRA into federated frameworks, respectively. More recently, methods like FFA-LoRA [42], FedEx-LoRA [41], Fed-SB [40], and RoLoRA [9] optimize LoRA in homogeneous-compute FL by addressing the inexactness in LoRA aggregation caused by separately averaging the B and A parameters. An orthogonal direction is explored by works such as HetLoRA [10], FlexLoRA [4], and FLoRA [46] that address computational heterogeneity in LoRA-based FL fine-tuning by allowing clients to train LoRA parameters with different ranks. However, an examination of cross-device fine-tuning remains limited in this context as methods like FLoRA have communication costs that scale linearly with the number of clients and communication rounds, making fine-tuning particularly difficult in large-scale FL systems. The goal of Ravan is to design a PEFT method for FL that addresses data and computational heterogeneity, while scaling effectively to cross-device settings with a large number of clients and communication rounds. In this way, we overcome shortcomings in prior works and enable resource‑aware fine‑tuning.